hi. im taking algebra 1, and im having trouble understanding how to convert units. in the problmes that i have to do, we are givin some units, but i have to figure out what goes in the question mark. 1100ft/1 second times 3600seconds/ 1 hour times ?/? = 750 miles/1 hour. Thats my equation, now i dont quite understand how the answer comes out to be 750 miles/1 hour. So if you could please help me it would be greatly apprecitated. Thanks a buch.
Answer by hyh_123(8)

your question requires you to convert ft/secs into miles/hr
remember 1 mile = 5280ft
1 hr =60*60=3600
the speed per hour =1100*3600 secs
=3960000
your answer will come in ft per secs but you have to give your answer in
miles per hr.
=3960000/5280
=750m/hr